OpenStudy (kiera_):

Which of the following correctly demonstrates the associative property? A. a+b+c+d = abcd B. (ab)+(c+d) = (ac)+(b+d) C. (a*b)*c = b*(c*a) D. none of the above

OpenStudy (kiera_):

C. (a*b)*c = b*(c*a)

Directrix (directrix):

The associative property for multiplication would say: (a*b)*c = a*(b*c) It is a grouping property and does not change the order of the variables. The a, b, and c appear in the same order on both sides of the equation. That is why I think this is not correct for this problem: C. (a*b)*c = b*(c*a) Look at the other options.